# Controlling Nonequilibrium Bose-Einstein Condensation with Engineered Environments

### v3

@@ Abstract
Out of thermal equilibrium, bosonic quantum systems can Bose-condense away from the ground state, featuring a macroscopic occupation of an excited state or even of multiple states in the so-called Bose-selection scenario. In previous work, a theory was developed that predicts, in which states a driven-dissipative ideal Bose gas condenses. Here, we address the inverse problem: Given a target state with desired condensate fractions in certain single-particle states, how can this configuration be achieved by tuning available control parameters? Which type of experimental setup allows for flexible condensation control? We solve these problems, on the one hand, by developing a theory to solve the inverse problem based on linear programming methods. On the other, we propose a Bose ‘condenser’, experimentally implementable in a superconducting circuit, where targeted Bose condensation into eigenstates of a chain of resonators is driven through the coupling to artificial quantum baths, realized via auxiliary two-level systems. We further discuss the engineering of transition points between different Bose condensation configurations, which may find application for amplification, heat-flow control, and the design of highly-structured quantum baths.
@@
Non-equilibrium Bose-Einstein condensation (BEC) has been widely explored in different platforms [1], such as photons in dye-filled cavities [2, 3, 4, 5, 6, 7, 8], excitons [9, 10, 11] and exciton-polaritons [12, 13, 14, 15, 16, 17, 18, 19] in (cavity) semiconductor heterostructures. BEC in these systems results from the interplay of thermalization with pump and loss, whose ratio determines the condensate mode. Non-equilibrium BEC of a different kind is predicted to occur, however, also at conserved particle number, when other mechanisms deprive the ground-state of its privileged role. Examples are open systems subject to time-periodic driving or to a strong competition between heating and cooling mechanisms [20, 21, 22, 23, 24]. Steady states with multiple condensates in Bose-selected modes arise here from the nonequilibrium quantum-jump kinetics. Controlling their (fragmented) condensation pattern on demand is an appealing perspective for the design of, e.g., quantum signal amplifiers, multimode emitters, structured artificial quantum baths or heat-transport regulators. However, while a theory explaining Bose selection (BS) has been developed [20, 21], approaches to systematically engineer nonequilibrium BS, turning it into a practical resource, are missing. The key challenges are (i) how to realize controllable quantum-jump networks in realistic experimental conditions and (ii) how to solve the inverse problem of finding values of the control knobs yielding a target BS pattern. In this work, we solve both problems by designing protocols to control non-equilibrium BEC of photons in a superconducting circuit using synthetic reservoirs given by driven-damped qubits.
@@
for the density operator $\hat{\rho}_{S}$, with dissipators $D[\hat{L}_{ij}]\hat{\rho}_{S}=\hat{L}_{ij}\hat{\rho}_{S}\hat{L}_{ij}^{\dagger}-\{\hat{L}_{ij}^{\dagger}\hat{L}_{ij},\hat{\rho}_{S}\}/2$, Hamiltonian $\hat{H}_{S}=\sum_{i}E_{i}\hat{n}_{i}$, rates $R_{ij}$ for quantum jumps from $j$ to $i$, jump operators $\hat{L}_{ij}=\hat{c}^{\dagger}_{i}\hat{c}_{j}$, and annihilation and number operators $\hat{c}_{i}$ and $\hat{n}_{i}=\hat{c}^{\dagger}_{i}\hat{c}_{i}$ for a boson in the single-particle eigenstate $i$ with energy $E_{i}$. The rates are assumed to realize a fully-connected network, implying a unique steady state [20], which becomes the equilibrium state, only when $R_{ji}/R_{ij}=\exp[-\beta(E_{i}-E_{j})]$, with inverse temperature $\beta$. Although the bosons are non-interacting, the coupling to the bath(s) makes the problem interacting, as the dissipator is quartic in the $\hat{c}_{i}$. As a result, the equation of motion for the mean occupations $n_{i}=\langle\hat{n}_{i}\rangle(t)=\mathrm{tr}[\hat{\rho}_{S}(t)\hat{n}_{i}]$ depends on two-particle correlations, $\dot{n}_{i}=\sum_{j}\big{[}R_{ij}\langle\hat{n}_{j}(1+\hat{n}_{i})\rangle-R_{ji}\langle\hat{n}_{i}(1+\hat{n}_{j})\rangle\big{]}$, as part of an infinite hierarchy of equations. This hierarchy can be truncated through a mean-field approximation $\langle\hat{n}_{i}\hat{n}_{j}\rangle\approx n_{i}n_{j}$, which yields nonlinear mean-field equations for the $n_{i}$ [20, 21]. The nonlinearity is related to Bose statistics, giving rise to a dependence of the many-body rate on the occupation a particle jumps to, known as bosonic enhancement (or stimulated emission). It is responsible for the emergence of BS in the steady state, satisfying $\dot{n}_{i}=0$, where a single or multiple selected states acquire a macroscopic occupation $\propto N$ in the large-$N$ limit, while the occupation of all other states saturates [sketch in Fig. 1(a)]. The set of selected states $S$ (and its complement $\bar{S}$) can be predicted by considering a large-$N$ expansion $n_{i}=\nu_{i}N+\sum_{\alpha=1}^{\infty}\nu_{i}^{(\alpha)}/N^{\alpha-1}$. It reveals that condensation is ruled, in leading order, by the rate-asymmetry matrix $A_{ij}=R_{ij}-R_{ji}$ through the set of (in)equalities [20, 21]
@@
which ensure that $n_{i}\geq 0$ for all $i$ [as reviewed in the Supplemental Material (SM) [25]]. The inequalities are understood elementwise and $\nu_{X}$ and $A_{XY}$ denote subvectors and matrix blocks, respectively, with $X,Y\in\{S,\bar{S}\}$.
@@
Bose condenser. We first propose a concrete realistic experimental implementation given by an array of $M$ microwave resonators in a superconducting circuit [28, 29, 30, 31, 32, 33, 34], each dispersively coupled to an ancillary transmon qubit subject to coherent microwave driving and loss, which represents a narrow-band artificial bath (hereafter described by a spin 1/2) [Fig. 1(c)]. Owing to the harmonicity of the resonator spectrum, the array hosts noninteracting microwave photons [33]. The dynamics is described by the master equation
@@
for the density matrix $\hat{\rho}$ of the combined system, where $\gamma_{\ell}$ is the decay rate of the $\ell$th spin and $\hat{\sigma}_{\ell}^{\alpha}$ ($\alpha=x,y,z,\pm$) are Pauli matrices. The Hamiltonian $\hat{H}_{S}$ of the bosonic system is given by $\hat{H}_{S}=\sum_{\ell}\omega_{\ell}\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}-\sum_{\ell,\ell^{\prime}}J_{\ell\ell^{\prime}}\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell^{\prime}}=\sum_{i}E_{i}\hat{c}_{i}^{\dagger}\hat{c}_{i},$ where $\hat{a}_{\ell}$ and $\hat{a}_{\ell}^{\dagger}$ are the annihilation and creation operators of a photon in the $\ell$th harmonic resonator, $\omega_{\ell}$ are the resonator transition frequencies and $J_{\ell\ell^{\prime}}>0$ are the tunnelling strengths. We choose the example of $\hat{H}_{S}$ describing a 1D chain, with $J_{\ell\ell^{\prime}}=J$ for nearest neighbours $\ell$ and $\ell^{\prime}$. For weak disorder, $|\omega_{\ell}-\omega_{\ell^{\prime}}|\sim J_{\ell\ell^{\prime}}$, the eigenmodes $i$ will be delocalized over several lattice sites. The artificial-bath Hamiltonian $\hat{H}_{B}=\sum_{\ell}\delta_{\ell}\left(\hat{\sigma}^{z}_{\ell}+r_{\ell}\hat{\sigma}^{x}_{\ell}\right)/2$ describes the transmon spins (in a frame rotating at the driving frequency), where $\delta_{\ell}$ is the detuning of the drive and $r_{\ell}$ is the ratio between its Rabi frequency and $\delta_{\ell}$. The coherent drive thus makes the spin precess on the Bloch sphere around a direction determined by the mixing angle $\theta_{\ell}=\arctan(r_{\ell})/2$. In the dispersive-coupling regime, the spins are far detuned from the resonators, such that resonant exchange of excitations is strongly unfavored, implying photon number conservation as desired (weak particle loss is discussed later), and effective photon-photon interactions induced by the coupling to the spins can also be neglected. The resonator-spin coupling is described by the Hamiltonian $\hat{H}_{SB}=\sum_{\ell}\chi_{\ell}\hat{a}^{\dagger}_{\ell}\hat{a}_{\ell}\hat{\sigma}_{\ell}^{z}$, provided the mean number of photons in a resonator does not exceed a critical value [35, 29]. For the parameters used below, tens to hundreds of photons per site are allowed, sufficient for BS. Condensation control will be achieved through the coherent drive and by tuning the value of the dispersive coupling $\chi_{\ell}$, in order to leverage the impact of each artificial bath on the quantum-jump network. Control of $\chi_{\ell}$ is implemented, for instance, with frequency-tuneable transmons, by adapting their detuning from the resonators [36, 37], or ones with tuneable coupling [38, 39]. Note that the use of nonlinear (two-level) elements as artificial bath, rather than additional cavity modes [40, 41, 42, 43], is crucial: the dispersive coupling to the latter would only give a mutual state-independent energy shift [29], rather than the density-density-like coupling in $\hat{H}_{SB}$.
@@
Engineered quantum-jump rates. Although $\hat{H}_{SB}$ commutes with the boson number $\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}$ in a single resonator, preventing particle loss, it has sizeable matrix elements between the non-site-local modes $i$, giving rise to non-zero off-diagonal elements of $R_{ij}$ tuneable via the spins. This is achieved by adapting the coherent-control parameters $\delta_{\ell}$ and $r_{\ell}$ such that the spin’s excitation energy, $\mathcal{E}_{\ell}=\delta_{\ell}\sqrt{1+r_{\ell}^{2}}$, matches the energy spacing between single-particle states |$\ket{i}$ and |$\ket{j}$ in the array. The coupling then drives a coherent excitation exchange involving a |$\ket{i}\leftrightarrow\ket{j}$ transition, with a matrix element $\chi_{ij}^{(\ell)}=r_{\ell}\chi_{\ell}M_{ij}^{(\ell)}/\sqrt{1+r_{\ell}^{2}}$ where $M_{ij}^{(\ell)}=\bra{i}\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\ket{j}$. In the meantime, the strong spin damping drags the spin to its drive-dependent steady state, interrupting the coherent flip-flop of the excitation and resulting in dissipative quantum jumps at rates $R_{ij}$ and $R_{ji}$. Here the spin relaxation time plays a role similar to a short bath correlation time in the more conventional Born-Markov scenario with a large bath. We derive the rates $R_{ij}=\sum_{\ell}R_{ij}^{(\ell)}$ from Eq. (S33) in the SM, where the spins are traced out in favour of an effective bath spectral density [25]. For the $\ell$th spin, tuned to resonate with the energy spacing $E_{ij}\equiv E_{i}-E_{j}>0$ and for $\delta_{\ell}>0$, we find
$R^{(\ell)}_{ij}=\gamma_{-}^{(\ell)}|\chi_{ij}^{(\ell)}|^{2}\mathcal{S}_{\ell}(E_{ij}),\ R^{(\ell)}_{ji}=\gamma_{+}^{(\ell)}|\chi_{ij}^{(\ell)}|^{2}\mathcal{S}_{\ell}(E_{ij}),$ (4)
@@
where $\mathcal{S}_{\ell}(\omega)=2\Gamma_{\ell}/[(\omega-|\mathcal{E}_{\ell}|)^{2}+\Gamma_{\ell}^{2}]$ is a Lorentzian quantum-noise spectrum of linewidth $\Gamma_{\ell}=\gamma_{\ell}[3-1/(1+r_{\ell}^{2})]/4$. For $\delta_{\ell}<0$, the rates have the same form (4), but with $\gamma^{(\ell)}_{\pm}$ exchanged. The effective decay constants $\gamma_{-}^{(\ell)}=\tan^{4}\theta_{\ell}/(1+\tan^{4}\theta_{\ell})$ and $\gamma_{+}^{(\ell)}=(1+\tan^{4}\theta_{\ell})^{-1}$ involve the spin’s mixing angle $\theta_{\ell}$. Hence, the direction of spin precession on the Bloch sphere determines the quantum-jump ratio $R^{(\ell)}_{ij}/R^{(\ell)}_{ji}=\gamma_{-}^{(\ell)}/\gamma_{+}^{(\ell)}$, which can be understood intuitively: For $r_{\ell}\ll 1$, the $\hat{\sigma}_{\ell}^{-}$ decay brings the spin close to its ground (excited) state for $\delta_{\ell}>0$ ($\delta_{\ell}<0$), which strongly favors processes that lower (raise) the energy in the system. For $r_{\ell}\gg 1$, $\hat{\sigma}_{\ell}^{-}$ brings the spin in a superposition of its two energy eigenstates, enabling both transitions lowering and raising the system energy. Intermediate values of $r_{\ell}$ are optimal to ensure a sizeable quantum-jump matrix element $\chi_{ij}^{(\ell)}\propto r_{\ell}$ while remaining in the weak-driving regime desired in experiments. Overall, each bath induces rates increasing or lowering the energy, which are enhanced around the peak of $\mathcal{S}_{\ell}$. Alternative control scenarios based on a time-periodic modulation of the bosonic system, would make the control more difficult, as they alter also the shape of the modes $i$, which become Floquet states, and give rise to side-band transitions (see, e.g., Refs. [20, 24]).
@@
The validity of Eq. (4) requires the hierarchy $E_{ij}\gg\gamma_{\ell}\gg|\chi_{ij}^{(\ell)}|$ between the system energy gap $E_{ij}$, the spin decay rates $\gamma_{\ell}$ and the transition matrix element $\chi_{ij}^{(\ell)}$. This ensures that off-resonant system-spin energy transfer mechanisms are negligible and that the spins behave as Markovian memoryless effective baths, only ‘filtering’ the noise spectrum of the environment. The above hierarchy defines the operating regime of the proposed Bose condenser, which is easily met in state-of-the-art superconducting devices [32]. Given that system gaps are of the order of the tunnelling strength $J$, we may choose, e.g., a typical value $J/2\pi\sim 30$ MHz and decay rates $\gamma_{\ell}\sim 1-10$ MHz (realized, e.g., in [32]). The values of the Rabi frequencies $|\delta_{\ell}|r_{\ell}$ used in the following are lower than $1.5J\sim 2\pi\times 40$ MHz, thus meeting standard values for microwave drives on transmon qubits. We then obtain $|\chi_{ij}^{(\ell)}|\sim 10^{-1}J$ by also restricting $\chi_{j}$ to a maximal value $\chi_{\rm max}\sim 0.15J\sim 2\pi\times 4.5$ MHz [28]. This in turn yields engineered rates $R_{ij}\sim 10^{-2}J$, much stronger than typical photon loss rates for microwave resonators in circuit QED [29] and guarantees that, even in case losses are not compensated for (as we propose below), the system forms a prethermal Bose-selected steady-state well before serious particle loss occurs.
@@
Figure 2: (a) Single-particle spectrum of the five-site chain. The arrow numbers indicate the reservoir spin tuned to induce the related transition. (b), (d)-(f) Controlled BS in states |$\ket{0}$, |$\ket{2}$, |$\ket{3}$: control parameters $\chi_{\ell}$ and asymmetry-matrix network (arrows and widths reflect $\mathrm{sign}(A_{ij})$ and $|A_{ij}|$, respectively) [(b)], single-particle dynamics [(d)], steady state [(e)] and mean-field dynamics for $N=50$ (solid) and $N=500$ (dashed) [(f)]. Symbols and line types are explained in the main text. (c) Controlled BEC into single modes |$\ket{1}$ and |$\ket{2}$, and corresponding values of $\chi_{\ell}$.
@@
Programmable BECs. To convert BS into a control problem, we identify controllable coefficients $\bm{z}$ leveraging the different bath contributions as $z_{\ell}=\mathrm{sign}(\delta_{\ell})(\chi_{\ell}/\chi_{\rm max})^{2}$, leading to a decomposition of the total asymmetry matrix as $A=\sum_{\ell}z_{\ell}A^{(\ell)}$. We then construct a recipe to reverse-engineer a $\bm{z}$ giving a target BS pattern $\bm{\nu}$. From the conditions (2) we derive a new set of inequalities for $\bm{z}$,

### v4

# Controlling Nonequilibrium Bose Condensation with Engineered Environments

@@ Abstract
Out of thermal equilibrium, bosonic quantum systems can Bose-condense away from the ground state, featuring a macroscopic occupation of an excited state or even of multiple states in the so-called Bose-selection scenario. In previous work, a theory was developed that predicts, in which states a driven-dissipative ideal Bose gas condenses. Here, we address the inverse problem: Given a target state with desired condensate fractions in certain single-particle states, how can this configuration be achieved by tuning available control parameters? Which type of experimental setup allows for flexible condensation control? We solve these problems, on the one hand, by proposing a Bose ‘condenser’, experimentally implementable in a superconducting circuit, where targeted Bose condensation into eigenstates of a chain of resonators is driven through the coupling to artificial quantum baths, realized via auxiliary two-level systems. On the other, we develop a theory to solve the inverse problem based on linear programming methods. We further discuss the engineering of transition points between different Bose condensation configurations, which may find application for amplification, heat-flow control, and the design of highly-structured quantum baths.
@@
Non-equilibrium Bose-Einstein condensation (BEC) has been widely explored in different platforms [1], such as photons in dye-filled cavities [2, 3, 4, 5, 6, 7, 8], excitons [9, 10, 11] and exciton-polaritons [12, 13, 14, 15, 16, 17, 18, 19] in (cavity) semiconductor heterostructures. BEC in these systems results from the interplay of thermalization with pump and loss, whose ratio determines the condensate mode. Non-equilibrium BEC of a different kind is predicted to occur, however, also at conserved particle number, when other mechanisms deprive the ground-state of its privileged role. Examples are open systems subject to time-periodic driving or to a strong competition between heating and cooling mechanisms [20, 21, 22, 23, 24]. Steady states with multiple condensates in Bose-selected modes arise here from the nonequilibrium quantum-jump kinetics. Controlling their (fragmented) condensation pattern on demand is an appealing perspective for the design of, e.g., quantum signal amplifiers, multimode emitters, structured artificial quantum baths or heat-transport regulators. However, while a theory explaining Bose selection (BS) has been developed [20, 21], approaches to systematically engineer nonequilibrium BS, turning it into a practical resource, are missing. The key challenges are (i) how to realize controllable quantum-jump networks in realistic experimental conditions and (ii) how to solve the inverse problem of finding values of the control knobs yielding a target BS pattern. In this work, we solve both problems by proposing a concrete experimental setup where nonequilibrium BEC of photons can be controlled in a superconducting circuit using synthetic reservoirs and by developing methods to reverse-engineer control parameters yielding desired condensation patterns. We demonstrate the success of this procedure in simulations with realistic values of experimental parameters and propose an application in the design of a quantum switch for heat transport.
@@
for the density operator $\hat{\rho}_{S}$. Given single-particle eigenstates |$\ket{i}$ with energy $E_{i}$, Eq. (1) involves the Hamiltonian $\hat{H}_{S}=\sum_{i}E_{i}\hat{n}_{i}$, dissipators $D[\hat{L}_{ij}]\hat{\rho}_{S}=\hat{L}_{ij}\hat{\rho}_{S}\hat{L}_{ij}^{\dagger}-\{\hat{L}_{ij}^{\dagger}\hat{L}_{ij},\hat{\rho}_{S}\}/2$ with rates $R_{ij}$ for quantum jumps from |$\ket{j}$ to |$\ket{i}$, jump operators $\hat{L}_{ij}=\hat{c}^{\dagger}_{i}\hat{c}_{j}$, and annihilation and number operators $\hat{c}_{i}$ and $\hat{n}_{i}=\hat{c}^{\dagger}_{i}\hat{c}_{i}$ for a boson in the $i$th single-particle state. The rates are assumed to realize a fully-connected network, implying a unique steady state [20], which becomes the equilibrium state, only when $R_{ji}/R_{ij}=\exp[-\beta(E_{i}-E_{j})]$, with inverse temperature $\beta$. Although the bosons are non-interacting, the coupling to the bath(s) makes the problem interacting, as the dissipator is quartic in the $\hat{c}_{i}$. The hierarchy of Eqs. for the $m$-point correlators $\mathrm{tr}[\hat{\rho}_{S}\hat{n}_{i_{1}}\hat{n}_{i_{2}}\dots\hat{n}_{i_{m}}]$, resulting from this interacting problem, can be truncated through a mean-field approximation, which yields nonlinear kinetic Eqs. of motion for the $n_{i}=\mathrm{tr}[\hat{\rho}_{S}\hat{n}_{i}]$, namely $\dot{n}_{i}=\sum_{j}\big{[}R_{ij}{n}_{j}(1+{n}_{i})-R_{ji}{n}_{i}(1+{n}_{j})\big{]}$  [20, 21] (see the SM [25] for a brief review). The nonlinearity is related to Bose statistics, giving rise to a dependence of the many-body rate on the occupation a particle jumps to, known as bosonic enhancement (or stimulated emission). It is responsible for the emergence of BS in the steady state, satisfying $\dot{n}_{i}=0$, where a single or multiple selected states acquire a macroscopic occupation $\propto N$ in the large-$N$ limit, while the occupation of all other states saturates [sketch in Fig. 1(a)]. The set of selected states $S$ (and its complement $\bar{S}$) can be predicted by considering a large-$N$ expansion $n_{i}=\nu_{i}N+\sum_{\alpha=1}^{\infty}\nu_{i}^{(\alpha)}/N^{\alpha-1}$. It reveals that condensation is ruled, in leading order, by the rate-asymmetry matrix $A_{ij}=R_{ij}-R_{ji}$ through the set of (in)equalities [20, 21]
@@
which ensure that $n_{i}\geq 0$ for all $i$. The inequalities are understood elementwise and $\nu_{X}$ and $A_{XY}$ denote subvectors and matrix blocks, respectively, with $X,Y\in\{S,\bar{S}\}$.
@@
A superconducting “Bose condenser”. We propose an experimental implementation given by an array of $M$ microwave resonators in a superconducting circuit [28, 29, 30, 31, 32, 33, 34]. Each resonator is dispersively coupled to an ancillary transmon qubit subject to coherent driving and loss, which implements a narrow-band artificial bath (hereafter described by a spin 1/2) [Fig. 1(c)]. Owing to the harmonicity of the resonator spectrum, the array hosts noninteracting microwave photons [33], whose condensation we aim to control. The dynamics is described by the master equation
@@
for the density matrix $\hat{\rho}$ of the combined resonators-spins system, where $\gamma_{\ell}$ is the decay rate of the $\ell$th spin and $\hat{\sigma}_{\ell}^{\alpha}$ ($\alpha=x,y,z,\pm$) are Pauli matrices. The Hamiltonian $\hat{H}_{S}$ of the bosonic system is given by $\hat{H}_{S}=\sum_{\ell}\omega_{\ell}\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}-\sum_{\ell,\ell^{\prime}}J_{\ell\ell^{\prime}}\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell^{\prime}}=\sum_{i}E_{i}\hat{c}_{i}^{\dagger}\hat{c}_{i},$ where $\hat{a}_{\ell}$ and $\hat{a}_{\ell}^{\dagger}$ are the annihilation and creation operators of a photon in the $\ell$th resonator, which has transition frequency $\omega_{\ell}$, and $J_{\ell\ell^{\prime}}>0$ are the tunnelling strengths. We choose the example of $\hat{H}_{S}$ describing a 1D chain, with $J_{\ell\ell^{\prime}}=J$ for nearest neighbours $\ell$ and $\ell^{\prime}$. We include a weak disorder, $|\omega_{\ell}-\omega_{\ell^{\prime}}|\sim J_{\ell\ell^{\prime}}$, to break symmetries inducing multiple identical level spacings in the spectrum. Concretely, we use $\omega_{\ell}/J=\ell/10+\varepsilon_{\ell}$, where $\varepsilon_{\ell}$ are random numbers uniformly distributed in $[0,1)$ [25], though the specific values are not important for the target physics. These values of $\omega_{\ell}$ are also chosen, such that the eigenmodes |$\ket{i}$ are still delocalized over several lattice sites for the system-sizes considered. These features will be useful for realizing efficient dissipation engineering. The artificial-bath Hamiltonian $\hat{H}_{B}=\sum_{\ell}\delta_{\ell}\left(\hat{\sigma}^{z}_{\ell}+r_{\ell}\hat{\sigma}^{x}_{\ell}\right)/2$ describes the transmon spins (in a frame rotating at the driving frequency), where $\delta_{\ell}$ is the detuning of the drive and $r_{\ell}$ is the ratio between its Rabi frequency and $\delta_{\ell}$. In the dispersive-coupling regime, the spins are far detuned from the resonators. Resonant exchange of excitations is thus strongly unfavored, implying photon number conservation as desired (weak particle loss is discussed later). Other system–artificial-bath couplings as used, e.g., in Refs. [32, 35] would involve particle exchange instead. Moreover, in the dispersive regime, undesired photon-photon interactions induced by the coupling to the spins can also be neglected. The resonator-spin coupling is described by the Hamiltonian $\hat{H}_{SB}=\sum_{\ell}\chi_{\ell}\hat{a}^{\dagger}_{\ell}\hat{a}_{\ell}\hat{\sigma}_{\ell}^{z}$, provided the mean number of photons in a resonator does not exceed a critical value [36, 29]. For the parameters used below, tens to hundreds of photons per site are allowed, sufficient for BS. Condensation control will be achieved through the coherent drive and by tuning the value of the dispersive coupling $\chi_{\ell}$, in order to leverage the impact of each artificial bath on the quantum-jump network. Control of $\chi_{\ell}$ is implemented, for instance, with frequency-tuneable transmons, by adapting their detuning from the resonators [37, 38], or ones with tuneable coupling [39, 40]. Note that the use of nonlinear (two-level) elements as artificial bath, rather than additional cavity modes [41, 42, 43, 44], is crucial: the dispersive coupling to the latter would only give a mutual state-independent energy shift [29], rather than the density-density-like coupling in $\hat{H}_{SB}$.
@@
Engineered quantum-jump rates. Although $\hat{H}_{SB}$ commutes with the boson number $\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}$ in a single resonator, preventing particle loss, it has sizeable matrix elements between the non-site-local modes |$\ket{i}$, giving rise to non-zero off-diagonal elements of $R_{ij}$ tuneable via the spins. Intuitively, this enables for dissipation engineering with the following mechanism. The coherent-control parameters $\delta_{\ell}$ and $r_{\ell}$ are adapted such that the spin’s excitation energy, $\mathcal{E}_{\ell}=\delta_{\ell}\sqrt{1+r_{\ell}^{2}}$, matches the energy spacing between states |$\ket{i}$ and |$\ket{j}$ in the array. The coupling drives a coherent excitation exchange involving a |$\ket{i}\leftrightarrow\ket{j}$ transition, with a matrix element $\chi_{ij}^{(\ell)}=r_{\ell}\chi_{\ell}M_{ij}^{(\ell)}/\sqrt{1+r_{\ell}^{2}}$, where $M_{ij}^{(\ell)}=\bra{i}\hat{a}_{\ell}^{\dagger}\hat{a}_{\ell}\ket{j}$, which we shall compute below. This excitation flip-flop is interrupted by the strong spin damping, which drags the spin to its drive-dependent steady state and results in dissipative quantum jumps in the bosonic chain at rates $R_{ij}$ and $R_{ji}$. Here the spin relaxation time plays a role similar to a short bath correlation time in the more conventional Born-Markov scenario with a large bath.
To derive the rates $R_{ij}=\sum_{\ell}R_{ij}^{(\ell)}$ of Eq. (1) from Eq. (S35), consider first a single reservoir-spin coupled to the $\ell$-th resonator (further technical details on the derivation are given in the SM [25]). The generalization to multiple reservoirs is straightforward, as their rates simply sum up. Representing Eq. (S35) in a diagonal basis for the spin, terms corresponding to non-secular (off-diagonal) elements of the Kossakowski matrix of the dissipator are off-resonant and can be neglected in rotating-wave approximation (RWA). This is justified provided the spin level splitting $|\mathcal{E}_{\ell}|$ is much larger than its decay rate $\gamma_{\ell}$, namely $|\mathcal{E}_{\ell}|\gg\gamma_{\ell}$ (Approximation I). Equation (S35) becomes
$\dot{\hat{\rho}}=-i\left[\sum_{i}E_{i}\hat{n}_{i}-\sum_{i,j}\chi_{ij}^{(\ell)}\hat{L}_{ij}\left(\frac{\hat{\sigma}_{\ell}^{z}}{r_{\ell}}+\hat{\sigma}_{\ell}^{x}\right)-\frac{\mathcal{E}_{\ell}}{2}\hat{\sigma}_{\ell}^{z},\hat{\rho}\right]\\ +\gamma_{\ell}^{+}D[\hat{\sigma}_{\ell}^{+}]\hat{\rho}+\gamma_{\ell}^{-}D[\hat{\sigma}_{\ell}^{-}]\hat{\rho}+\gamma_{\ell}^{z}D[\hat{\sigma}_{\ell}^{z}]\hat{\rho}\ ,$ (4)
including $\mathcal{E}_{\ell}=\delta_{\ell}\sqrt{1+r_{\ell}^{2}}$ and the dressed decay rates $\gamma_{\ell}^{+}=\gamma_{\ell}\cos(\theta_{\ell})^{4},\ \gamma_{\ell}^{-}=\gamma_{\ell}\sin(\theta_{\ell})^{4},\ \gamma_{\ell}^{z}=\gamma_{\ell}r_{\ell}^{2}/[4(1+r_{\ell}^{2})]$, where $\theta_{\ell}=\arctan(r_{\ell})/2$ is the spin’s mixing angle.
[…]
$\mathcal{S}_{\ell}^{\pm}(\omega)=\frac{\gamma_{\ell}^{\mp}}{\gamma_{\ell}^{+}+\gamma_{\ell}^{-}}\frac{2\Gamma_{\ell}}{(\omega\pm\mathcal{E}_{\ell})^{2}+\Gamma_{\ell}^{2}}\ ,$ (5)
characterized by a linewidth $\Gamma_{\ell}=(\gamma_{\ell}/4)\left[3-1/(1+r_{\ell}^{2})\right]$. For a positive $E_{ij}>0$ and detuning $\delta_{\ell}>0$, the rates can be rewritten as
$R^{(\ell)}_{ij}=\gamma_{\ell}^{-}|\chi_{ij}^{(\ell)}|^{2}\mathcal{S}_{\ell}(E_{ij}),\ R^{(\ell)}_{ji}=\gamma_{\ell}^{+}|\chi_{ij}^{(\ell)}|^{2}\mathcal{S}_{\ell}(E_{ij}),$ (6)
@@
with $\mathcal{S}_{\ell}(\omega)=2\Gamma_{\ell}/[(\omega-|\mathcal{E}_{\ell}|)^{2}+\Gamma_{\ell}^{2}]$. For $\delta_{\ell}<0$, the rates have the same form (6), but with $\gamma_{\ell}^{\pm}$ exchanged. The driven spin dynamics determines the quantum-jump ratio $R^{(\ell)}_{ij}/R^{(\ell)}_{ji}=\gamma_{\ell}^{-}/\gamma_{\ell}^{+}=\tan(\theta_{\ell})^{4}$ via the mixing angle $\theta_{\ell}$. For $\delta_{\ell}>0$ and $\theta_{\ell}\ll 1$, the spin dissipator in Eq. 4 drags the spin to its ground state, favouring the processes $\hat{L}_{ij}\hat{\sigma}^{+}_{\ell}$ which increase the system energy, but blocking the inverse processes $\hat{L}_{ij}^{\dagger}\hat{\sigma}^{-}_{\ell}$ which lower the energy. The opposite occurs for $\theta_{\ell}\approx\pi/2$. Intermediate values of $\theta_{\ell}$ allow both processes with a finite rate and are optimal to ensure a sizeable matrix element $\chi_{ij}^{(\ell)}\propto r_{\ell}$, while complying with the weak-driving regime desired in experiments. Summarizing, each artificial bath can be tuned to increase or decrease the system energy with a controllable rate ratio and the corresponding rates are enhanced around the peak of $\mathcal{S}_{\ell}(\omega)$. Alternative control scenarios based on a Floquet modulation of the bosonic system may be possible, but would make the control more difficult. Indeed, strong driving alters the shape of the modes |$\ket{i}$, which become Floquet states, and gives rise also to side-band quantum jumps (see, e.g., Refs. [20, 24]).
@@
The validity of Eq. (6) requires the hierarchy $E_{ij}\gg\gamma_{\ell}\gg|\chi_{ij}^{(\ell)}|$ between the system energy gap $E_{ij}$, the spin decay rates $\gamma_{\ell}$ and the transition matrix elements $\chi_{ij}^{(\ell)}$, resulting from Approximations I-III. This hierarchy defines the operating regime of the proposed Bose condenser, which is easily met in state-of-the-art superconducting devices [32]. In particular, we envision implementations involving a number of resonators of the order of ten, each coupled to a superconducting qubit, which are close to the setup of recent experiments [32, 31] and suffice for the potential applications described below. Given that system gaps are of the order of the tunnelling strength $J$ for such system sizes, we may choose a value $J/2\pi\sim 30$ MHz and decay rates $\gamma_{\ell}\sim 1-10$ MHz (realized, e.g., in [32]). The Rabi frequencies $|\delta_{\ell}|r_{\ell}$ used in our examples are lower than $1.5J\sim 2\pi\times 40$ MHz in value, thus meeting standard constraints for microwave drives on transmons. We then obtain $|\chi_{ij}^{(\ell)}|\sim 10^{-1}J$ by also restricting $\chi_{j}$ to a maximal value $\chi_{\rm max}\sim 0.15J\sim 2\pi\times 4.5$ MHz [28]. In turn, this yields engineered quantum-jump rates $R_{ij}\sim 10^{-2}J$, much stronger than typical photon loss rates for microwave resonators in circuit QED [29]. Even in case losses are not compensated for (as we propose below), the system can thus form a Bose-selected steady-state well before serious particle loss occurs.
@@
Figure 2: (a) Single-particle spectrum of the five-site chain. The arrow numbers indicate the reservoir-spin used to induce the dissipative transition between the corresponding linked states. (b), (d), (e), (f) show the controlled BS in states |$\ket{0}$, |$\ket{2}$, |$\ket{3}$, namely the values of the control parameters $\chi_{\ell}$ and the resulting asymmetry-matrix network [(b)– the arrows and their thickness represent $\mathrm{sign}(A_{ij})$ and $|A_{ij}|$, respectively], the single-particle dynamics [(d)], the steady state as a function of the total particle number $N$ [(e)] and the mean-field dynamics for $N=50$ (solid) and $N=500$ (dashed) [(f)]. Symbols and line types are explained in the main text. (c) Controlled BEC into single modes |$\ket{1}$ and |$\ket{2}$, and corresponding values of $\chi_{\ell}$.
@@
Programmable BECs. To convert BS into a control problem in the Bose condenser, we identify controllable coefficients $\bm{z}$ leveraging the different bath contributions as $z_{\ell}=\mathrm{sign}(\delta_{\ell})(\chi_{\ell}/\chi_{\rm max})^{2}$, leading to a decomposition of the total asymmetry matrix as $A=\sum_{\ell}z_{\ell}A^{(\ell)}$. We then construct a recipe to reverse-engineer a $\bm{z}$ giving a target BS pattern $\bm{\nu}$. From the conditions (2) we derive a new set of inequalities for $\bm{z}$,